# Distance-Preserving Spatial Representations in Genomic Data

### v2

@@ Appendix E Additional Experiment Results
Evaluation Protocals. Moran’s I [28] is a correlation coefficient that measures the overall spatial autocorrelation of a dataset. Intuitively, for a given gene, it measures how one spot is similar to others surrounding it. If the spots are attracted (or repelled) by each other, it implies they are not independent. Thus, the presence of autocorrelation indicates the spatial pattern of gene expression. Moran’s I value ranges from –1 to 1, where a value close to 1 indicates a clear spatial pattern, a value close to 0 indicates random spatial expression, and a value close to –1 indicates a chessboard-like pattern. To evaluate the spatial variability of a given gene, we calculate the Moran’s I using the following formula,
@@
The experiment results are shown in Figure 12, Figure 13, Table I. Figure 12 shows the gene expression levels of six marker genes extracted by VAE. For each gene in each panel, the left subplot shows the gene expression levels in the spatial coordinates, and the right subplot shows the 2D-UMAP of the latent representation of the corresponding model. The cells are dispersed in the latent space. The olfactory bulb region, with genes Calb2, Gng4, S100a5, and Febp7 enriched, is well separated from other regions. Table I shows that enforcing distance-preserving property induces stronger spatial autocorrelations on out-of-sample datasets. Figure 13 shows that the proposed method can improve isometric property of the learned latent space.
Figure 13: Visualization of latent representation space obtained from scVI (left) and scVI regularized with distance-preserving loss (right). More isotropic and homogeneous ellipses indicate more distance-preserving [56].
@@
Appendix F Additional Background of VAE
CVAE. A conditional variational encoder (CVAE) [43, 52] is an extension of VAE, where an additional conditional input $y\in\mathcal{Y}$ can be specified in the encoder and decoder network
@@ Appendix F Additional Background of VAE
$\beta$-VAE. $\beta$-VAE [19] incorporates a temporature term of the KL loss term in the VAE model:
$-\beta\cdot D_{\text{KL}}(q(z|x)||p_{\theta}(z))+\mathbb{E}_{z\sim q(\cdot|x)}\left[\log p_{\theta}(x|z)\right].$
@@
scVI (single-cell Variational Inference). scVI [33] is a leading deep generative model for analyzing single-cell RNA sequencing (scRNA-seq) data. It employs a CVAE architecture to model gene expression patterns across individual cells. In scVI, the input consists of gene expression counts for each cell, while the model aims to learn a latent representation that captures biological variability. The model accounts for technical factors such as library size and batch effects, which are treated as observed variables. Specifically, scVI uses the normalized gene expression counts as the target variable, with the library size included as a nuisance variable to be inferred. Batch information, when available, is incorporated as conditional information to help disentangle technical from biological variability. This approach allows scVI to perform various downstream tasks, including normalization, batch correction, imputation, and dimensionality reduction, making it a versatile tool for single-cell transcriptomics analysis.
Markovian Hierarchical VAE. A Markovian Hierarchical Variational Autoencoder (MHVAE) [24, 44] is a generalization of a VAE that extends to multiple hierarchies over latent variables. Under this formulation, latent variables are interpreted as generated from other higher-level, more abstract patterns. Let $T>1$ denote the total number of hierarchical levels; the objective of an MHVAE is similar to VAE, except that there are now $T$ latent variables whose joint distribution can be decomposed into Markov transitions:
[…]
@@
Deriving for the ELBO of the MHVAE can follow a similar procedure as of VAE:
@@
However, a key caveat is that the expectation over two random variables $\{z_{t-1},z_{t+1}\}$ for every timestep, the variance of its Monte Carlo estimate could potentially be higher than a term that is estimated using only one random variable per timestep. As it is computed by summing up $T-1$ consistency terms, the final estimated value of the ELBO may have a high variance for large $T$ values. Instead, we can derive an equivalent version that does not pose this issue:
@@
The changed last term is called denoising matching term following the variational diffusion model literature [34]. We learn desired denoising transition step $p_{\theta}(z_{t-1}|z_{t})$ as an approximation to tractable, variational denoising transition step $q_{\phi}(z_{t-1}|z_{t},x)$.

### v3

@@ Appendix E Additional Experiment Details
Evaluation Protocals. Moran’s I [li2007beyond] is a correlation coefficient that measures the overall spatial autocorrelation of a dataset. Intuitively, for a given gene, it measures how one spot is similar to others surrounding it. If the spots are attracted (or repelled) by each other, it implies they are not independent. Thus, the presence of autocorrelation indicates the spatial pattern of gene expression. Moran’s I value ranges from –1 to 1, where a value close to 1 indicates a clear spatial pattern, a value close to 0 indicates random spatial expression, and a value close to –1 indicates a chessboard-like pattern. To evaluate the spatial variability of a given gene, we calculate Moran’s I using the following formula,
@@
The experiment results are shown in Figure 12, Figure 14, Table I. Figure 12 shows the gene expression levels of six marker genes extracted by VAE. For each gene in each panel, the left subplot shows the gene expression levels in the spatial coordinates, and the right subplot shows the 2D-UMAP of the latent representation of the corresponding model. The cells are dispersed in the latent space. The olfactory bulb region, with genes Calb2, Gng4, S100a5, and Febp7 enriched, is well separated from other regions. Table I shows that enforcing the distance-preserving property induces stronger spatial autocorrelations on out-of-sample datasets. Figure 14 shows that the proposed method can improve the isometric property of the learned latent space.
Figure 14: Visualization of latent representation space obtained from scVI (left) and scVI regularized with distance-preserving loss (right). More isotropic and homogeneous ellipses indicate more distance-preserving [yonghyeon2021regularized].
@@
Transfer Learning Setup. Across all of our experiments, we used highly variable genes (HVGs), meaning that they were selected based on variance in gene expression across cells. We chose HVGs over spatially variable genes (SVGs) because HVGs capture the most informative expression patterns for representation learning, and they are more likely to be consistently expressed across different datasets and brain regions.
The adequate expression of these selected genes across both training and test datasets is ensured by the design of our experiment, where we intentionally used “similar” datasets as training and testing pairs. Specifically, the four selected datasets – abbreviated as VMBSP, VMBSPS2, VMBSA, and VMBSAS2 – all come from the V1 mouse brain sagittal area. This similarity ensures that the overall gene expression patterns are also highly similar across the four datasets, including those of the highly variable genes. In fact, a total of $74\%$ out of the top 100 HVGs that were selected appear across all four datasets, with the proportion increasing further when only considering pairwise overlaps, as shown in Table II:
TABLE II: Pairwise overlap ratio of shared genes across datasets
Dataset VMBSP VMBSPS2 VMBSA VMBSAS2 VMBSP $100\%$ $96\%$ $77\%$ $79\%$ VMBSPS2 $96\%$ $100\%$ $78\%$ $81\%$ VMBSA $77\%$ $78\%$ $100\%$ $92\%$ VMBSAS2 $79\%$ $81\%$ $92\%$ $100\%$
Finally, the top 100 HVGs that are shared across the four datasets are selected, and we aligned them in the same order to form our dataset vectors. Therefore, this procedure ensures HVGs are adequately expressed in our experiments.
Appendix F Spatial Reconstruction Error
This section describes the detailed procedure for computing the spatial reconstruction error defined in Section 5.1.
Recall that the spatial reconstruction error is defined as solving the following optimization problem:
Here $\textbf{A}\in\mathbb{R}^{m,m}$ is an orthogonal matrix, $b\in\mathbb{R}^{m}$, and $c$ is a scalar. The objective value is equivalent ot the Procrustes distance between datasets $\mathbf{D}_{1},\mathbf{D}_{2}\in\mathbb{R}^{n\times m}$. The optimization problem is solved mainly via singular value decomposition (SVD). Specifically:
  1. 1.
    Both the rescaling term $c$ and translation term $b$ are determined through the normalization, where the former is set to the ratio of their Forbenius norm $\|\mathbf{D}_{2}\|_{F}/\|\mathbf{D}\|_{f}$, and the latter is set to the difference of their centroid $(\mathbf{1}^{\top}\mathbf{D}_{2}-\mathbf{1}^{\top}\mathbf{D}_{1})/n$.
  2. 2.
    Then, the rotation matrix $\mathbf{A}$ is computed via SVD of the cross-covariance matrix of the normalized data $\tilde{\mathbf{D}}_{1}:=c\cdot\mathbf{D}_{1}+b$ and $\mathbf{D}_{2}$, finding the optimal orthogonal transformation matrix that “aligns” the two datasets the best to minimize their Frobenius distance: (i) The cross-covariance matrix is computed as $\Sigma=\tilde{\mathbf{D}}_{1}^{\top}\mathbf{D}_{2}$; (ii) Compute its SVD $\Sigma=U\Lambda V^{\top}$; (iii) Set $\mathbf{A}=UV^{\top}$.
In our implementation, we directly use the SciPy Python package [2020SciPy-NMeth], where $\mathbf{A}$, $b$, and $c$ are all computed implicitly within intermediate procedures.
Appendix G Additional Background of VAE
CVAE. A conditional variational encoder (CVAE) [sohn2015learning, wu2023counterfactual] is an extension of VAE, where an additional conditional input $y\in\mathcal{Y}$ can be specified in the encoder and decoder network
@@ Appendix G Additional Background of VAE
$\beta$-VAE. $\beta$-VAE [higgins2017beta] incorporates a temperature term of the KL loss term in the VAE model:
$\beta\cdot D_{\text{KL}}(q(z|x)||p_{\theta}(z))+\mathbb{E}_{z\sim q(\cdot|x)}\left[\log p_{\theta}(x|z)\right].$
@@
scVI (single-cell Variational Inference). scVI [lopez2018deep] is a leading deep generative model for analyzing single-cell RNA sequencing (scRNA-seq) data. It employs a CVAE architecture to model gene expression patterns across individual cells. In scVI, the input consists of gene expression counts for each cell, while the model aims to learn a latent representation that captures biological variability. The model accounts for technical factors such as library size and batch effects, which are treated as observed variables. Specifically, scVI uses the normalized gene expression counts as the target variable, with the library size included as a nuisance variable to be inferred. Batch information, when available, is incorporated as conditional information to help disentangle technical from biological variability. This approach allows scVI to perform various downstream tasks, including normalization, batch correction, imputation, and dimensionality reduction, making it a versatile tool for single-cell transcriptomics analysis.
Markovian Hierarchical VAE. A Markovian Hierarchical Variational Autoencoder (MHVAE) [kingma2016improved, sonderby2016ladder] is a generalization of a VAE that extends to multiple hierarchies over latent variables. Under this formulation, latent variables are interpreted as generated from other higher-level, more abstract patterns. Let $T>1$ denote the total number of hierarchical levels; the objective of an MHVAE is similar to VAE, except that there are now $T$ latent variables whose joint distribution can be decomposed into Markov transitions:
$p_{\theta}(x)$ $=\int p_{\theta}(x|z_{1:T})dz_{1:T}$ (19)
@@
Deriving the ELBO of the MHVAE can follow a similar procedure to that of VAE:
@@
However, a key caveat is that the expectation over two random variables $\{z_{t-1},z_{t+1}\}$ for every timestep, the variance of its Monte Carlo estimate could potentially be higher than a term that is estimated using only one random variable per timestep. As it is computed by summing up T1T-1 consistency terms, the final estimated value of the ELBO may have a high variance for large $T$ values. Instead, we can derive an equivalent version that does not pose this issue:
@@
The modified last term is called denoising matching term following the variational diffusion model literature [luo2022understanding]. We learn desired denoising transition step $p_{\theta}(z_{t-1}|z_{t})$ as an approximation to tractable, variational denoising transition step $q_{\phi}(z_{t-1}|z_{t},x)$.